Find the electric field due to a 2 C charge at a distance 1cm apart

Grade:12th pass

1 Answers

Deepak Sharma
5 years ago
We know that,
    E=kq/r^2=9×10^9×2/(0.1)^2=1.8×10^12N/C.
So the electric field due to the charge 2c at a distance0.1m is 1.8×10^12N/C